OpenStudy (anonymous):

Given the function f(x) = 4(x + 1)2 - 3, indicate the shifts that will affect the location of the vertex, and explain what effect they will have. Use complete sentences. f(x-2) f(x) - 2 f(2x) 2*f(x)

OpenStudy (anonymous):

@calculusfunctions

OpenStudy (anonymous):

explain please

OpenStudy (calculusfunctions):

@1017squad do you want @DanJS to just do the work for you and give you the answers or do you want someone to teach you? As a teacher I'm here to teach. It is very unsettling to see everyone simply showing off and just giving out the solutions. How is one supposed to learn?!

OpenStudy (anonymous):

well to be honest im not really trying to learn this but it wont hurt and i understand what your saying

OpenStudy (anonymous):

@calculusfunctions i need an explanation on why these work so id rather learn it

OpenStudy (calculusfunctions):

OK then, given a quadratic function in vertex form\[y =a(x -p)^{^{2}}+q\]If a > 1 then the parent function is vertically stretched by a factor of a, If a < 1 then the parent function is vertically compressed by a factor of a, If p < 0 (x + p) then the parent function is shifted p units left, If p > 0 (x - p) then the parent function is shifted p units right, If q < 0 then the parent function is shifted q units down, If q > 0 then the parent function is shifted q units up The parent function referred to above is\[f(x) =x ^{2}\]

OpenStudy (anonymous):

im honestly lost do you think you can draw it out?

OpenStudy (calculusfunctions):

Can you make the connections? In your question, do know a, p, and q?

OpenStudy (anonymous):

no sir

OpenStudy (calculusfunctions):

Also one thing I forgot to add. If a < 0 (a is negative), then the parabola's direction of opening is downward.

OpenStudy (anonymous):

ok i still dont get it

OpenStudy (calculusfunctions):

OK so you have\[y =4(x +1)^{2}-3\]and in general\[y =a(x -p)^{2}+q\]Do you not see what a, p, and q values are in your question?

OpenStudy (anonymous):

i see it in the question now

OpenStudy (anonymous):

ok so whats next

OpenStudy (calculusfunctions):

So a = ?, p = ? and q = ? First tell me these values.

OpenStudy (anonymous):

i see the numbers but how do i get the values

OpenStudy (calculusfunctions):

What do you mean? The corresponding numbers are the values.

OpenStudy (anonymous):

so they are 4,1,and 3

OpenStudy (calculusfunctions):

Almost! a = 4 p = -1 q = -3 Do you see that?

OpenStudy (calculusfunctions):

Now identify the transformations based on the description I gave you for each parameter.

OpenStudy (anonymous):

compressed, shifted to the right and down?

OpenStudy (anonymous):

i mean left

OpenStudy (calculusfunctions):

For example, if we had\[y =5(x +2)^{2}+7\]Then a = 5 indicates a vertical stretch (since a > 0) by a factor of 5, p = -2 indicates a shift of 2 units left (because p < 0), q = 7 indicates a shift of 7 units up (since q > 0)

OpenStudy (anonymous):

so how do i apply this to my problem?

OpenStudy (calculusfunctions):

Exactly the same way using your values of a, p, and q and the explanation I gave you.

OpenStudy (calculusfunctions):

First you have a = 4, correct? Then from the notes I gave you, what does it mean?

OpenStudy (anonymous):

so i do that to get one or two of those 4 choices?

OpenStudy (anonymous):

yes it means the parent function is vertically compressed

OpenStudy (calculusfunctions):

No, a = 4 > 1, which means it is vertically STRETCHED by a factor of 4. You're not even paying attention!! I didn't write a lesson for my own amusement!

OpenStudy (anonymous):

sorry ok so whats next?

OpenStudy (calculusfunctions):

Now you have p = -1. Please read carefully about what I said about p and write your answer.

OpenStudy (anonymous):

then the parent function is shifted p units left,

OpenStudy (calculusfunctions):

Yes but in your question that would be 1 unit to the left. Now finally what about q?

OpenStudy (anonymous):

then the parent function is shifted q units down

OpenStudy (anonymous):

-3 units down right

OpenStudy (calculusfunctions):

YES!

OpenStudy (calculusfunctions):

Do you get it now? I would probably save those notes if I were you, for future reference.

OpenStudy (calculusfunctions):

Alright then Good Luck!!! I'll be signing out for now.

OpenStudy (anonymous):

no one more question please you could just give me an aswer and an explanation